use the variable expression 3x-1.
evaluate the expression when x= -3,-2,-1,0,1,2, and 3.
THANK YOU VERY MUCH I REALLY DO APPRECIATE IT.

Answer by AnlytcPhil(1810)

use the variable expression 3x-1.
evaluate the expression when x= -3,-2,-1,0,1,2, and 3. 
THANK YOU VERY MUCH I REALLY DO APPRECIATE IT. 

1. Write the expression 

3x - 1

Now put (-3) where x is

3(-3) - 1

Then finish this out: 3(-3) is -9, so write -9 in place of 3(-3)

-9 - 1

Then finish this and get -10.

So when you evaluate the expression 3x - 1 for x = -3, you get -10.

---------------------------------------------------------------------

2. Write the expression 

3x - 1

Now put (-2) where x is

3(-2) - 1

Then finish this out: 3(-2) is -6, so write -6 in place of 3(-2)

-6 - 1

Then finish this and get -7.

So when you evaluate the expression 3x - 1 for x = -2, you get -7.

---------------------------------------------------------------------

3. Write the expression 

3x - 1

Now put (-1) where x is

3(-1) - 1

Then finish this out: 3(-1) is -3, so write -3 in place of 3(-1)

-3 - 1

Then finish this and get -4.

So when you evaluate the expression 3x - 1 for x = -1, you get -4.

---------------------------------------------------------------------

4. Write the expression 

3x - 1

Now put (0 where x is

3(0 - 1

Then finish this out: 3(0)is 0 so write 0 in place of 3(0)

0 - 1

Then finish this and get -1.

So when you evaluate the expression 3x - 1 for x = 0, you get -1.

---------------------------------------------------------------------

5. Write the expression 

3x - 1

Now put (1) where x is

3(1) - 1

Then finish this out: 3(1) is 3, so write 3 in place of 3(1)

3 - 1

Then finish this and get 2.

So when you evaluate the expression 3x - 1 for x = 1, you get 2.

---------------------------------------------------------------------

6. Write the expression 

3x - 1

Now put (2) where x is

3(2) - 1

Then finish this out: 3(2) is 5, so write 6 in place of 3(2)

6 - 1

Then finish this and get 5.

So when you evaluate the expression 3x - 1 for x = 2, you get 5.

---------------------------------------------------------------------

7. Write the expression 

3x - 1

Now put (3) where x is

3(3) - 1

Then finish this out: 3(3) is 9, so write 9 in place of 3(3)

9 - 1

Then finish this and get 8.

So when you evaluate the expression 3x - 1 for x = 3, you get 8.

---------------------------------------------------------------------

Now put the results of the seven steps above in a table, like this

x  |  3x-1
----------
-3 |  -10
-2 |   -7
-1 |   -4
 0 |   -1
 1 |    2
 2 |    5
 3 |    8

Edwin